OpenStudy (anonymous):

solve for the variable. x/3 = 8/12

OpenStudy (anonymous):

x = 2

OpenStudy (blank ):

(x)/(3)=(8)/(12) Multiply each term in the equation by 3. x=2

OpenStudy (anonymous):

I asked these questions before then I forgot them shay lol my freaken school network shut off!

OpenStudy (anonymous):

what about w/15 = 3/5 ?

OpenStudy (blank ):

w = 9

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks :)

OpenStudy (blank ):

ǝɯoɔlǝʍ ǝɹ,noʎ :)

OpenStudy (anonymous):

i just told you those answers lol you have bad memory

OpenStudy (anonymous):

I know lol not my fault I got a sucky computer idk whats happening man

OpenStudy (anonymous):

lol it's cool